OpenStudy (electronia):

Simplify 4x + 8y - 3x + 2y

pooja195 (pooja195):

Hi there Welcome to openstudy Start by combining like terms! 3x-3x=? 8y+2y= ?

OpenStudy (electronia):

6x and 10y?

pooja195 (pooja195):

I meant 4x-3x

OpenStudy (electronia):

7x

pooja195 (pooja195):

10y is correct

pooja195 (pooja195):

4-3=? we are subtracting

OpenStudy (electronia):

Oh! Ok. 1x

OpenStudy (anonymous):

the answer is x+10y

OpenStudy (electronia):

Oh! Thank you!

pooja195 (pooja195):

ok good so we combine those and get 1x+10y :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (electronia):

Yay! Thank you! :)

pooja195 (pooja195):

:)

OpenStudy (electronia):

How do you factor an equation?